A small ball rolls horizontally off the edge of a tabletop that is 1,2 m high. It strikes the floor at a point 1,52 m horizontally away from the edge of the table.

1. How long is the ball in the air?
2. What is the speed at the instant it leaves the table?
1
Expert's answer
2021-08-25T12:25:33-0400

1)

"t=\sqrt{\frac{2h}g}=0.5~s,"

2)

"v_y=\sqrt{2gh}=4.85~\frac ms,"

"v_x=l/t=3.04~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=5.72~\frac ms."